# Atomic Calibration of LLMs in Long-Form Generations

### v1

@@ 5.2 Results
Overall, the tested LLMs are not well-calibrated at the atomic fact level. Table 1 lists our main atomic calibration results, which indicate that different confidence elicitation methods yield varying calibration scores. Although there is no universally accepted threshold for low ECE, a well-calibrated model typically achieves an ECE close to 1%, as shown in (Guo et al., 2017) and (Zhu et al., 2023). However, even with the most robust method, Gen-Binary, the ECE scores remain around 10%, indicating a significant calibration gap. Among the models, Qwen2-7B-Instruct demonstrates slightly better calibration compared to the other two.
@@
Confidence fusion can further improve both atomic and macro calibration. Table 4 presents the results of various confidence fusion strategies at the atomic level (with response-level results in Appendix D). The best results are consistently achieved by fusion methods. Among these strategies, the weighted average (WAvg) is the most effective. The benefits of the fusion strategies are further discussed in §6.1. Interestingly, we find that combining methods within the same confidence type (such as Dis-Rating with Dis-Context) does not lead to improved calibration (in Appendix D). A case study showing the effectiveness of confidence fusion is in Figure 4.
@@
Table 3: Atomic Calibration Results. All the numbers are in percentages. Our two key findings are: (1) for generative methods, there is no significant difference in calibration between larger and smaller models; (2) however, with discriminative methods, larger models tend to provide better calibration.
@@
Table 4: Atomic calibration results of confidence fusion strategies for Llama3-8B-Instruct, demonstrating the effectiveness of confidence fusion in combining generative and discriminative confidence estimates.
@@ 6.1 Confidence Methods Alignment
(a) Atomic Level Confidences.
(b) Response Level Confidences.
Figure 2: Heatmaps comparing the Spearman correlation between generative and discriminative confidence elicitation methods in Llama3-8B-Instruct on WildHallu. Warmer colors are for higher correlations.
@@
To further explore the reasons behind the improvements provided by confidence fusion, we show the correlation between different confidence elicitation methods in Figure 2 (using WildHallu as the study case and more results are in Appendix E). Our findings are summarized as follows:
Confidence methods within the same type are better aligned. In Figure 2, warmer colors indicate higher Spearman correlation scores. Confidence elicitaton methods of the same type (top left for generative and bottom right for discriminative) show stronger correlations compared to those across different types. This helps to explain why fusing generative and discriminative methods are effective, since these two types capture different aspects of uncertainty and are complementary to each other.
@@ 6.2 Confidence Across Different Positions
7 Related Work
Atomic Facts Generation and Verification. Long-form responses often contain both correct and incorrect statements, which impact the overall factuality assessments. Min et al. (2023) propose breaking long responses into atomic facts and calculating the precision of these fact pieces to determine the overall factuality score. Wei et al. (2024) and Zhao et al. (2024) extend this paradigm by expanding the dataset to include more domains beyond biographies. Song et al. (2024) design VeriScore for diverse long-form generation tasks that feature both verifiable and unverifiable content. Chiang and Lee (2024) introduce D-FActScore, specifically designed for content with ambiguous entities. Overall, the approach of breaking long-form responses into atomic facts for verification is nowadays widely accepted in factuality assessment.
Uncertainty and Calibration in Long-form Generations. Existing research on uncertainty estimation and calibration primarily focuses on multiple-choice or short-form questions (Zhu et al., 2023; Kuhn et al., 2022; Lin et al., 2023; Tian et al., 2023; Ulmer et al., 2024). However, limited work has explored calibration for long-form generations. Huang et al. (2024) proposed a unified calibration framework for all text generation tasks, comparing distributions of both correctness and the associated confidence of responses. Band et al. (2024) introduced linguistic calibration, where models explicitly express their uncertainty during long-form generation. Zhang et al. (2024) proposed Luq, an uncertainty estimation method tailored to long-form generation, demonstrating its effectiveness in ensembling different LLMs. None of the aforementioned studies have systematically examined calibration at a fine-grained level. Our work aims to fill this gap with a focus on factuality.
@@ 8 Conclusion
We introduce atomic calibration, a fine-grained approach for evaluating LLM calibration at the atomic claim level, addressing the limitations of traditional response-level calibration in long-form generation. Our experiments show that atomic calibration is well-suited for long-form generation and complements macro calibration. We also demonstrated that combining different confidence elicitation methods leads to better calibration results and atomic calibration can provide deeper insights into model confidence during the generation process. This approach offers a more reliable way to assess and improve the trustworthiness of LLM outputs in real-world applications.
@@ Appendix A Atomic Calibration Metrics
where $o_{i}$ is the true fraction of positive instances in bin $i$, $e_{i}$ is the mean of the post-calibrated probabilities for the instances in bin $i$, and $P(i)$ is the empirical probability (fraction) of all instances that fall into bin $i$. The lower the $ECE$ value is, the better a model is calibrated.
@@
The Brier score measures the accuracy of probabilistic predictions. For unidimensional predictions, it is strictly equivalent to the mean squared error as applied to predicted probabilities. Suppose that on each of the $n$ occasions an event can occur in only one of $r$ possible classes or categories and on one such occasion $i$, the forecast probabilities are $f_{i1},f_{i2}$, $\ldots f_{ir}$, that the event will occur in classes $1,2,\ldots r$, respectively. The $r$ classes are chosen to be mutually exclusive and exhaustive so that
$\sum_{i=1}^{r}f_{ij}=1,i=1,2,3,\ldots n.$
The Brier Score $P$ is then defined by
$P=\frac{1}{n}\sum_{j=1}^{r}\sum_{i=1}^{n}\left(f_{ij}-E_{ij}\right)^{2},$
where $E_{ij}$ takes the value 1 or 0 according to whether the event occurred in class $j$ or not.
@@
Appendix B Prompts
Table 6: Prompt for Gen-Binary.
Table 7: Prompt for Gen-Multi.
Table 8: Prompt for Dis-Single.
Table 9: Prompt for Dis-Context.
Table 10: Prompt for Dis-Rating.
Appendix C Experiment Details
@@
Appendix D Confidence Fusion Results
Bios LongFact WildHallu
SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$
Gen-Binary 79.8 5.6 5.6 52.7 3.0 3.0 70.0 7.2 7.8
Dis-Rating 73.8 25.7 25.7 34.1 2.8 2.9 71.7 3.8 3.6
Dis-Context 55.4 33.2 34.0 29.7 5.4 5.6 65.9 9.1 9.5
[…]
HMean 81.7 11.5 11.5 54.5 2.0 2.0 78.7 3.2 3.0
ProdConf 82.2 7.7 7.1 54.9 14.2 14.3 78.7 12.2 12.1
WAvg 81.8 10.9 11.0 55.6 1.5 0.9 77.2 5.0 5.0
Table 11: Macro calibration results of confidence fusion strategies for Llama3-8B-Instruct.
Bios LongFact WildHallu
SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$
Gen-Binary 74.9 8.0 8.5 64.1 2.5 2.5 73.9 10.3 10.3
Dis-Rating 55.0 41.5 41.4 40.8 4.2 4.4 60.4 16.7 16.9
Dis-Context 79.7 8.2 8.3 47.9 4.2 4.1 72.3 5.7 6.1
[…]
HMean 83.5 5.8 6.2 62.5 1.9 2.1 77.9 6.5 6.6
ProdConf 83.8 13.1 13.0 62.5 10.6 10.4 78.0 7.2 7.1
WAvg 82.0 8.0 8.0 65.1 1.8 2.0 78.4 8.9 8.9
Table 12: Macro calibration results of confidence fusion strategies for Mistral-7B-Instruct.
Bios LongFact WildHallu
SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$
Gen-Binary 72.4 5.7 5.4 67.6 1.8 2.0 72.2 6.4 6.5
Dis-Rating 63.0 41.0 40.7 29.9 4.4 4.4 54.0 8.9 9.1
Dis-Context 66.5 14.4 14.8 40.6 3.9 3.9 66.8 3.7 4.0
[…]
HMean 76.2 7.8 8.0 62.4 1.4 1.4 74.7 2.3 2.2
ProdConf 76.0 12.9 12.6 62.5 10.4 10.4 74.9 12.7 12.7
WAvg 75.7 7.5 8.4 66.4 1.2 1.4 75.6 4.5 4.3
Table 13: Macro calibration results of confidence fusion strategies for Qwen2-7B-Instruct.
@@ Appendix D Confidence Fusion Results
Table 14: Atomic calibration results of confidence fusion strategies for Mistral-7B-Instruct.
@@
Table 15: Atomic calibration results of confidence fusion strategies for Qwen2-7B-Instruct.
Appendix E Confidence Alignment

### v2

@@ 5.2 Results
Overall, the tested LLMs are not well-calibrated at the atomic fact level. Table 1 lists our main atomic calibration results. Although there is no universally accepted threshold for low ECE, a well-calibrated model typically achieves an ECE close to 1%, as shown in (Guo et al., 2017) and (Zhu et al., 2023). However, even with the most robust method, Gen-Binary, the ECE scores remain around 10%, indicating a significant calibration gap. Among the models, Qwen2-7B-Instruct demonstrates slightly better calibration compared to the other two.
Models that appear well-calibrated at the response level still perform poorly at the atomic level
Figure 2 compares atomic and response-level scores for ECE and Brier Score across different datasets and confidence types. The data points consistently lie above the identity line, indicating that atomic-level errors are higher than response-level errors. This suggests that atomic calibration is crucial for fine-grained evaluation.
@@
The confidence fusion method considering confidence agreement outperforms other methods. Table 4 presents the results of various confidence fusion strategies at the atomic level (more results in Appendix F). The best performance is consistently achieved by AdjustedAlpha and DampedFusion. Notably, we observe that combining methods of the same confidence type (e.g., Dis-Rating with Dis-Context) does not lead to improved calibration. A case study demonstrating the effectiveness of confidence fusion is shown in Figure 11.
@@
Table 3: Atomic calibration results with different model sizes. All the numbers are in percentages.
@@
[…]
DampedFusion 5.0 15.6 84.7 3.5 9.8 80.0 4.8 12.4 87.9
Table 4: Atomic calibration results of different confidence fusion strategies for Llama3-8B-Instruct. The fusion results are based on Gen-Binary and Dis-Rating.
@@ 6.1 Confidence Methods Alignment
Figure 3: Heatmaps of Spearman Correlation between different confidences in Llama3-8B-Instruct on WildHallu. Warmer colors indicate higher correlations. Atomic level: left; response level: right.
To further explore the reasons behind the improvements provided by confidence fusion, we show the correlation between different confidence elicitation methods in Figure 3 (using WildHallu as the study case and more results are in Appendix G). Our findings are summarized as follows:
Confidence methods within the same type are better aligned. In Figure 3, warmer colors indicate higher Spearman Correlation scores. Confidence elicitaton methods of the same type (top left for generative and bottom right for discriminative) show stronger correlations compared to those across different types. This helps to explain why cross-category fusion strategies are effective, since these two types capture different aspects of uncertainty and are complementary to each other.
@@ 6.2 Confidence Across Different Positions
6.3 The Utilities of Atomic Calibration
While the primary goal of atomic calibration is to provide fine-grained calibration evaluation for models, we also explore its utilities in several downstream tasks, including: (1) Selective Question Answering (Kamath et al., 2020; Cole et al., 2023; Yang et al., 2023), which involves setting a confidence threshold to selectively reject low-confidence answers, ensuring that only high-confidence responses are retained; (2) LLM-Ensemble (Zhang et al., 2024), which leverages multiple models to generate responses to the same question, selecting the answer with the highest confidence, thereby combining the strengths of each model; and (3) Atomic Claims Reunion (Thirukovalluru et al., 2024; Jiang et al., 2024), which involves sampling multiple responses, breaking then into atomic claims, evaluating their confidence, and reassembling only high-confidence claims to produce a more reliable final answer. Among these applications, we observe consistent improvements in factuality with atomic-level examination. Detailed experimental settings and results can be found in Appendix C.
It is important to note that, unlike previous work on Selective Question Answering (Huang et al., 2024) and LLM-Ensemble (Zhang et al., 2024) for long-form generation, which mainly rely on atomic-level confidence estimation to enhance the overall quality of responses (with responses either being entirely accepted or rejected), Atomic Claims Reunion does not require an overall response-level score. Instead, it relies entirely on the confidence of atomic claims to select and combine the most accurate claims. This means the final answer may contain claims from different sampled answers. More importantly, we observe that models with better atomic-level calibration (e.g., Qwen2 in Table 7, Appendix C) exhibit greater improvements after the reunion process, emphasizing the importance of examining and refining atomic calibration.
@@ 7 Conclusion
Our main contributions are three-fold: (1) We systematically study atomic calibration, which evaluates confidence calibration at the level of individual atomic claims. Our experiments reveal that models that appear well-calibrated at the response level perform poorly at the atomic level. (2) To analyze confidence elicitation methods, we categorize them into discriminative and generative methods. We also propose two novel fusion strategies to combine the confidence scores based on confidence agreement. (3) Our atomic-level analysis provides further insights into confidence methods alignment and confidence changes during generation. We find with discriminative methods, models show decreasing confidence in atomic facts as generation progresses. In contrast, generative methods show the lowest average confidence in the middle of the generation. Last but not least, we demonstrate the utilities of atomic calibration and propose for future research on more fine-grained confidence in long-form generation.
@@
Third, in this work, we mainly focus on exploring different confidence elicitation methods and therefore do not apply post-hoc calibration techniques such as histogram binning or temperature scaling. Applying these methods makes it difficult to disentangle improvements due purely to elicitation from those due to recalibration. To isolate the contribution of our elicitation designs and to avoid conflating them with downstream post-processing effects, we report raw atomic and macro confidence scores, leaving a systematic study of post-hoc techniques to future work.
@@ Appendix A Atomic Calibration Metrics
where $o_{i}$ is the true fraction of positive instances in bin $i$, $e_{i}$ is the mean of the post-calibrated probabilities for the instances in bin $i$, and $P(i)$ is the empirical probability (fraction) of all instances that fall into bin $i$. The lower the $ECE$ value, the better a model is calibrated.
When labels are continuous values between 0 and 1, the ECE formulation can be generalized. Instead of binning instances based on binary outcomes, the continuous predictions are grouped into bins according to their predicted probability values. Specifically, the observed calibration error $o_{i}$ in each bin is the average of the continuous label values for the instances in that bin, and $e_{i}$ is the mean predicted probability for those instances. This ensures that the calibration error accounts for all possible real-valued outcomes within the range [0,1], providing a more nuanced measure of calibration when the labels are continuous.
@@
The Brier score measures the accuracy of probabilistic predictions. In binary classification, it compares the predicted probability of the positive class with the actual binary outcome (0 or 1). The Brier score is defined as:
$P=\frac{1}{n}\sum_{i=1}^{n}\left(\hat{y}_{i}-y_{i}\right)^{2},$
where $\hat{y}_{i}$ is the predicted probability for instance $i$ and $y_{i}\in\{0,1\}$ is the actual binary outcome.
For continuous labels in the range $[0,1]$, the Brier score can still be used, where $y_{i}$ is now a continuous value between 0 and 1. In this case, the Brier score becomes equivalent to the mean squared error (MSE) between predicted probabilities and the true values, and minimizing the Brier score for continuous labels is analogous to minimizing MSE. Both metrics aim to reduce the squared differences between predicted and true values, with lower scores indicating better calibration and accuracy.
@@
Spearman Correlation
Following Zhang et al. (2024), we calculate Spearman Correlation to assess whether samples with higher factuality have correspondingly higher confidence scores. Compared to Pearson Correlation, it focuses on assessing the rank correlation, is robust to outliers and does not require that data is in normal distribution.
Appendix B Statistics in Atomic and Macro Calibration
To assess the confidence of a model, we generate responses using various questions (e.g., $N$ questions). For each response, a single confidence score is too coarse-grained. Instead, we evaluate the confidence of each atomic claim, with an average of $M$ atomic claims per response. These individual confidences are then aggregated into a response-level confidence score.
Atomic calibration is computed over $MN$ data points, where $M$ is the average number of atomic claims per response, and $N$ is the number of responses. In contrast, response-level calibration is based on $N$ data points. This distinction highlights the trustworthiness of the model’s confidence at both the atomic and response levels, providing a more granular view of its performance.
From the above discussion, it follows that to ensure sufficient data points for atomic calibration, $MN$ must be large. In our datasets, $N$ typically exceeds 1,000, ensuring that $MN$ remains robust even when some responses have only a few calims.
The detailed generation statistics are further illustrated in Figures 5, 6, and 7. Figure 5 presents the average answer length, while Figure 6 shows the average number of atomic claims per answer. Finally, Figure 7 highlights the percentage of answers containing fewer than 10 atomic facts.
Figure 5: Average answer length (in words) for different models on Bios, longfact, and wildhallu.
Figure 6: Average number of atomic claims per answer for different models on Bios, longfact, and wildhallu.
Figure 7: Answer with less than 10 atomic facts (%) by dataset and model. Notably, Mistral-8$\times$7B-Instruct has 6% short answers. Human evaluation reveals that these responses are primarily instances where the model refuses to answer.
Appendix C Applications
Selective Question Answering:
Selective question answering involves setting a confidence threshold, which can be derived from a validation set, to selectively reject questions with low confidence. This approach aims to improve the overall factuality of the responses by eliminating potentially unreliable answers.
Using the Bios dataset, we evaluate the performance of three models: Llama3-8B-Instruct, Mistral-7B-Instruct, and Qwen2-7B-Instruct with Dis-Gen and Semantic Entropy (SE). We observe an improvement in overall factuality as we gradually rejected more questions (from 0% to 10%). The table below illustrates this trend, highlighting the utility of Dis-Gen in identifying accurate responses and improving selective question answering. Our comparison between Dis-Gen and Semantic Entropy (SE) indicates that Dis-Gen brings more significant improvements in factuality, suggesting that better calibration methods can substantially enhance the results of selective question answering.
LLM Ensemble:
In the LLM Ensemble method, we use three models to generate answers to the same question and select the response from the model with the highest confidence. This approach aims to enhance factuality by leveraging the strengths of each model. The Answer Distribution (AD) shows the proportion of the final response contributed by each model, highlighting the benefit of ensemble methods. The table below presents the results of applying this method on the Bios and WildHallu datasets, comparing two different selection strategies: DIS-Gen and SE.
The results demonstrate that the ensemble method with DIS-Gen significantly improves factuality. For instance, in the Bios dataset, the factuality score increases from 0.475 to 0.556, and in the WildHallu dataset, it increases from 0.655 to 0.752. In contrast, using SE results in no improvements, with factuality scores even lower than the best individual model (0.484 vs 0.502 and 0.671 vs 0.701) for the Bios and WildHallu datasets, respectively. These findings suggest that ensembling does not always guarantee better results, and the selection strategy, such as DIS-Gen, plays a crucial role in improving factuality.
Atomic Reunion:
In Atomic Reunion, for each question, we begin by sampling the model’s output five times (this is also what we need to calculate Gen-Dis. These outputs are then broken down into atomic claims, which are individual, verifiable statements. Each claim is evaluated for confidence, and only those with a high confidence level are retained. Subsequently, we prompt a LLM, such as GPT-4o, to reassemble the selected atomic claims into a cohesive and factually accurate response.
This method seeks to enhance factuality by utilizing smaller, more manageable pieces of information, allowing the LLM to generate a more reliable answer by combining only high-confidence atomic claims. The table below presents the factuality scores of the new answers generated through the Atomic Reunion approach. As observed, this approach leads to a significant improvement in factuality compared to the baseline models.
Refuse Rate Llama3-8B-Instruct Mistral-7B-Instruct Qwen2-7B-Instruct
SE
0% 0.475 0.403 0.502
5% 0.479 0.407 0.506
7.5% 0.483 0.411 0.511
10% 0.485 0.416 0.518
DIS-Gen
0% 0.475 0.403 0.502
5% 0.496 0.419 0.517
7.5% 0.511 0.438 0.533
10% 0.528 0.465 0.557
Table 5: Factuality Scores with Varying Refuse Rates for Selective Question Answering
Model Bios WildHallu
Factuality Scores Answer Distribution Factuality Scores Answer Distribution
DIS-Gen
Llama3-8B-Instruct 0.475 31% 0.655 29%
Mistral-7B-Instruct 0.403 23% 0.631 27%
Qwen2-7B-Instruct 0.502 46% 0.701 44%
Ensemble 0.556 / 0.752 /
SE
Llama3-8B-Instruct 0.475 18% 0.655 15%
Mistral-7B-Instruct 0.403 37% 0.631 44%
Qwen2-7B-Instruct 0.502 45% 0.701 41%
Ensemble 0.484 / 0.671 /
Table 6: LLM Ensemble Results on Bios and WildHallu Datasets
Model Before Atomic Reunion After Atomic Reunion
DIS-Gen
Llama3-8B-Instruct 0.475 0.501
Mistral-7B-Instruct 0.403 0.441
Qwen2-7B-Instruct 0.502 0.575
Table 7: Factuality Scores Before and After Atomic Reunion using DIS-Gen
Appendix D Reliability of Atomic Facts Generation and Verification
The processes of atomic fact generation and verification have been extensively studied and validated in prior work Min et al. (2023); Wei et al. (2024); Zhao et al. (2024). For instance, FActScore Min et al. (2023) reports an error rate of 2%. In this work, we leverage their pipeline while employing stronger models from GPT-3.5 and GPT-4o, to further enhance performance.
We, the authors, conducted additional tests comparing GPT’s atomic decompositions with ground-truth manual segmentations. We manually selected 30 samples for this evaluation. The results of our assessment are as follows:
  • Consistency: Over 10 trials, GPT demonstrated a high inter-run consistency of 95%, indicating stable and repeatable outcomes.
  • Error Rate: The error rate, which includes missing or overly segmented claims and misclassification of factuality, was measured at 6.4%. This error rate is manageable within the context of our calibration framework, suggesting that the model’s atomic fact generation is reliable.
Appendix E Experiment Details
@@
Appendix F Confidence Fusion Results
@@
[…]
DampedFusion 10.2 16.5 84.6 5.9 10.9 81.9 7.1 15.4 83.8
Table 8: Atomic calibration results of confidence fusion strategies for Mistral-7B-Instruct. The fusion results are based on Gen-Binary and Dis-Context.
@@
[…]
DampedFusion 10.2 15.8 84.5 2.6 9.3 82.9 5.2 13.3 84.4
Table 9: Atomic calibration results of confidence fusion strategies for Qwen2-7B-Instruct. The fusion results are based on Gen-Binary and Dis-Context.
Appendix G Confidence Alignment
@@
Appendix H Case Study
Appendix I Prompts
Table 11: Prompt for Gen-Binary.
Table 12: Prompt for Gen-Multi.
Table 13: Prompt for Dis-Single.
Table 14: Prompt for Dis-Context.
Table 15: Prompt for Dis-Rating.